# Onsager-Machlup Functional and Large Deviation Principle for Stochastic Hamiltonian Systems

### v1

@@ 3. Onsager-Machlup functional for stochastic Hamiltonian systems
(3.8) $\limsup\limits_{\epsilon\to 0}\mathbb{E}\left({\rm exp}\left\{cB_{3}\right\}\big{|}\left\|W^{\sigma}\right\|<\epsilon\right)=1$
@@
For the fourth term $B_{4}$, employing the same proof technique as for the third term $B_{3}$, we have
(3.9) $\limsup\limits_{\epsilon\to 0}\mathbb{E}\left({\rm exp}\left\{cB_{4}\right\}\big{|}\left\|W^{\sigma}\right\|<\epsilon\right)=1$
@@
$B_{5}$ $=\int_{0}^{1}\left\langle\sigma_{q}^{-2}(s)\left(\frac{\partial H}{\partial y_{p}}(y_{q},y_{p})-\frac{\partial H}{\partial\varphi_{p}}(\varphi_{q},\varphi_{p})\right),\dot{\varphi}_{q}(s)\right\rangle\,{\rm d}s$
$\leq C\left|\frac{\partial H}{\partial y_{p}}(y_{q},y_{p})-\frac{\partial H}{\partial\varphi_{p}}(\varphi_{q},\varphi_{p})\right|$
$\leq CL\left\|W^{\sigma}\right\|.$
Thus,
@@
(3.10) $\limsup\limits_{\epsilon\to 0}\mathbb{E}\left({\rm exp}\left\{cB_{5}\right\}\big{|}\left\|W^{\sigma}\right\|<\epsilon\right)=1$
for all $c\in\mathbb{R}$.
For the sixth term $B_{6}$, employing the same proof technique as for the fifth term $B_{5}$, we have
(3.11) $\limsup\limits_{\epsilon\to 0}\mathbb{E}\left({\rm exp}\left\{cB_{6}\right\}\big{|}\left\|W^{\sigma}\right\|<\epsilon\right)=1$
for all $c\in\mathbb{R}$.
For the first term $B_{1}$, in order to write it as a whole, we define
$\sigma^{-1}(t)$ $:=\begin{bmatrix}\sigma_{q}^{-1}(t)&0\\ 0&\sigma_{p}^{-1}(t)\\ \end{bmatrix},$ $\quad H^{\prime}(y):=\begin{pmatrix}\frac{\partial H}{\partial y_{p}}(y_{q},y_{p})\\ -\frac{\partial H}{\partial y_{q}}(y_{q},y_{p})\\ \end{pmatrix}$
$\quad W(t)$ $:=\begin{pmatrix}W_{q}(t)\\ W_{p}(t)\\ \end{pmatrix},$ $\,{\rm d}W(t):=\begin{pmatrix}\,{\rm d}W_{q}(t)\\ \,{\rm d}W_{p}(t)\\ \end{pmatrix}.$
Under the assumption of small perturbations, it is feasible to apply a Taylor series expansion to $H^{\prime}(y)$. Specifically, we have
$\quad H^{\prime}(y)$
$=\begin{pmatrix}\frac{\partial H}{\partial\varphi_{p}}(\varphi_{q},\varphi_{p})\\ -\frac{\partial H}{\partial\varphi_{q}}(\varphi_{q},\varphi_{p})\\ \end{pmatrix}+\begin{bmatrix}\frac{\partial^{2}H}{\partial\varphi_{q}\partial\varphi_{p}}(\varphi_{q},\varphi_{p})&\frac{\partial^{2}H}{\partial^{2}\varphi_{p}}(\varphi_{q},\varphi_{p})\\ -\frac{\partial^{2}H}{\partial^{2}\varphi_{q}}(\varphi_{q},\varphi_{p})&-\frac{\partial^{2}H}{\partial^{2}\varphi_{p}\partial\varphi_{q}}(\varphi_{q},\varphi_{p})\\ \end{bmatrix}\begin{pmatrix}W^{\sigma}_{q}(t)\\ W^{\sigma}_{p}(t)\\ \end{pmatrix}+\begin{pmatrix}R_{q}(t)\\ R_{p}(t)\\ \end{pmatrix}$
$:=H^{\prime}(\varphi)+J(H)W^{\sigma}+R(t).$
According to the properties of the Taylor expansion, when $H\in C^{3}_{b}(\mathbb{R}^{n}\times\mathbb{R}^{n},\mathbb{R})$ and $\left\|W^{\sigma}\right\|\leq\epsilon$, we can estimate the remainder term $R(t)$ as follows:
$\sup\limits_{t\in[0,1]}\left|R(t)\right|\leq k\epsilon^{2}.$
Hence, $B_{1}$ can be written as:
$B_{1}$ $=\int_{0}^{1}\left\langle\sigma_{q}^{-1}(s)\frac{\partial H}{\partial y_{p}}(y_{q},y_{p}),\,{\rm d}W_{q}(s)\right\rangle-\int_{0}^{1}\left\langle\sigma_{p}^{-1}(s)\frac{\partial H}{\partial y_{q}}(y_{q},y_{p}),\,{\rm d}W_{p}(s)\right\rangle$
$=\int_{0}^{1}\left\langle\sigma^{-1}(s)H^{\prime}(y),\,{\rm d}W(s)\right\rangle$
$=\int_{0}^{1}\left\langle\sigma^{-1}(s)H^{\prime}(\varphi),\,{\rm d}W(s)\right\rangle+\int_{0}^{1}\left\langle\sigma^{-1}(s)J(H)W^{\sigma},\,{\rm d}W(s)\right\rangle$
$\quad+\int_{0}^{1}\left\langle\sigma^{-1}(s)R(s),\,{\rm d}W(s)\right\rangle$
$:=B_{11}+B_{12}+B_{13}.$
The term $B_{11}$ has the same expression as $B_{2}$:
$B_{11}=\int_{0}^{1}\left\langle\sigma^{-1}(s)H^{\prime}(\varphi),\,{\rm d}W(s)\right\rangle.$
Due to $H\in C^{3}_{b}(\mathbb{R}^{n}\times\mathbb{R}^{n},\mathbb{R})$, we can show that $\sigma^{-1}(s)H^{\prime}(\varphi)\in L^{2}([0,1],\mathbb{R}^{n}\times\mathbb{R}^{n})$. Using the same method as item $B_{2}$ yields
(3.12) $\limsup\limits_{\epsilon\to 0}\mathbb{E}\left({\rm exp}\left\{cB_{11}\right\}\big{|}\left\|W^{\sigma}\right\|<\epsilon\right)=1$
for all $c\in\mathbb{R}$. In order to apply Lemma 2.10, we will express the term $B_{12}$ as a double stochastic integral with respect to $W$. We have
$B_{12}$ $=\int_{0}^{1}\left\langle\sigma^{-1}(s)J(H)W^{\sigma},\,{\rm d}W(s)\right\rangle$
$=\int_{0}^{1}\int_{0}^{s}\left\langle\sigma^{-1}(s)J(H)\sigma(t),\,{\rm d}W(t)\right\rangle\,{\rm d}W(s)$
$=\int_{0}^{1}\int_{0}^{1}\sigma^{-1}(s)J(H)\sigma(t)1_{t\leq s}\,{\rm d}W(t)\,{\rm d}W(s),$
where $1_{t\leq s}$ is an indicator function. Define
$F(s,t):=\sigma^{-1}(s)J(H)\sigma(t)1_{t\leq s}.$
Hence, $B_{12}=I_{2}(\tilde{F})$, where $\tilde{F}:=\frac{1}{2}(F+F^{*})$ is the symmetrization of $F$. According to Conditions $(C2)$ and $(C3)$, the operator $K(\tilde{F})$ is nuclear, and its trace can be computed as follows:
$Tr.\tilde{F}=Tr.{F}=\frac{1}{2}\int_{0}^{1}{{F}(t,t)}\,{\rm d}t=\frac{1}{2}\int_{0}^{1}Tr\left(J(H)\right)\,{\rm d}t.$
In addition, due to $H\in C^{2}$, we have
@@
$Tr\left(\frac{\partial^{2}H}{\partial\varphi_{q}\partial\varphi_{p}}(\varphi_{q},\varphi_{p})\right)=Tr\left(\frac{\partial^{2}H}{\partial\varphi_{p}\partial\varphi_{q}}(\varphi_{q},\varphi_{p})\right),$
and since the Hamiltonian equations possess a symplectic structure, we obtain:
$Tr\left(J(H)\right)=Tr\left(\frac{\partial^{2}H}{\partial\varphi_{q}\partial\varphi_{p}}(\varphi_{q},\varphi_{p})\right)-Tr\left(\frac{\partial^{2}H}{\partial\varphi_{p}\partial\varphi_{q}}(\varphi_{q},\varphi_{p})\right)=0.$
By Lemma 2.10 and Lemma 2.13, we have
(3.13) $\limsup\limits_{\epsilon\to 0}\mathbb{E}\left({\rm exp}\left\{cB_{12}\right\}\big{|}\left\|W^{\sigma}\right\|<\epsilon\right)=1$
for all $c\in\mathbb{R}$. Finally, we study the behaviour of the term $B_{13}$. For any $c\in\mathbb{R}$ and $\delta>0$, we have
(3.14) $\quad\mathbb{E}\left({\rm exp}\left\{cB_{13}\right\}\big{|}\left\|W^{\sigma}\right\|\leq\epsilon\right)$
$=\int_{0}^{\infty}{e^{x}\mathbb{P}\left(\left|c\int_{0}^{1}\left\langle\sigma^{-1}(s)R(s),\,{\rm d}W(s)\right\rangle\right|>x\big{|}\left\|W^{\sigma}\right\|\leq\epsilon\right)}\,{\rm d}x$
$\leq e^{\delta}+\int_{\delta}^{\infty}{e^{x}\mathbb{P}\left(\left|c\int_{0}^{1}\left\langle\sigma^{-1}(s)R(s),\,{\rm d}W(s)\right\rangle\right|>x\big{|}\left\|W^{\sigma}\right\|\leq\epsilon\right)}\,{\rm d}x.$
Define the martingale $M_{t}=c\int_{0}^{t}\langle\sigma^{-1}(s)R(s),\,{\rm d}W(s)\rangle$. We have the estimate about its quadratic variation
$\langle M_{t}\rangle=c^{2}\int_{0}^{t}{\left\|\sigma^{-1}(s)R(s)\right\|^{2}}\,{\rm d}s\leq C\epsilon^{4}$
$\mathbb{P}\left(\left|c\int_{0}^{1}\left\langle\sigma^{-1}(s)R(s),\,{\rm d}W(s)\right\rangle\right|>x,\left\|W^{\sigma}\right\|\leq\epsilon\right)\leq{\rm exp}\left\{-\frac{x^{2}}{2c\epsilon^{4}}\right\}.$
Then, by Lemma 2.11, we have
$\quad\mathbb{P}\left(\left|c\int_{0}^{1}\left\langle\sigma^{-1}(s)R(s),\,{\rm d}W(s)\right\rangle\right|>x\big{|}\left\|W^{\sigma}\right\|\leq\epsilon\right)$
$\leq{\rm exp}\left\{-\frac{x^{2}}{2c\epsilon^{4}}\right\}{\rm exp}\left\{c\left(\frac{\epsilon}{M}\right)^{-\frac{2}{1-2\alpha}}\right\}.$

### v2

@@ 3 Onsager-Machlup functional for stochastic Hamiltonian systems
$=\int_{0}^{1}\mathrm{Tr}\big{(}\sigma^{-1}(s,\varphi_{s})\left[JD_{y}\nabla H(\varphi_{s})-\left(D_{y}\sigma(s,\varphi_{s})\right)\sigma^{-1}(s,\varphi_{s})\times\left(J\nabla H(\varphi_{s})-\dot{\varphi}_{s}\right)\right]\sigma(s,\varphi_{s})\big{)}\,{\rm d}s$
$=\int_{0}^{1}\mathrm{Tr}\left(JD_{y}\nabla H(\varphi_{s})-\left(D_{y}\sigma(s,\varphi_{s})\right)\sigma^{-1}(s,\varphi_{s})\left(J\nabla H(\varphi_{s})-\dot{\varphi}_{s}\right)\right)\,{\rm d}s.$
Here, we used the cyclic property of the matrix trace: for any three matrices $A$, $B$, and $C$ of compatible dimensions, the identity $\mathrm{Tr}(ABC)=\mathrm{Tr}(BCA)=\mathrm{Tr}(CAB)$ holds. In particular, this allows us to cyclically rearrange the terms inside the trace when computing $\mathrm{Tr}\left(\bar{F}(s,s)\right)$. Additionally, due to the linearity of the trace, $\mathrm{Tr}(A+B)=\mathrm{Tr}(A)+\mathrm{Tr}(B)$, we can analyze $\mathrm{Tr}\left(JD_{y}\nabla H(\varphi_{s})\right)$ separately. Since $J$ is the standard symplectic matrix and $D_{y}\nabla H(\varphi_{s})$ represents the Hessian matrix of $H$, we can prove that $\mathrm{Tr}\left(JD_{y}\nabla H(\varphi_{s})\right)=0$.
By Lemma 2.21, we have
@@
$\quad\limsup\limits_{\epsilon\to 0}\mathbb{E}\left({\rm exp}\left\{cB_{\cdot 21}\right\}\big{|}\|W^{\sigma}\|<\epsilon\right)$ (15)
$={\rm exp}\left\{\frac{1}{2}\int_{0}^{1}\mathrm{Tr}\left(\left(D_{y}\sigma(s,\varphi_{s})\right)\sigma^{-1}(s,\varphi_{s})\left(J\nabla H(\varphi_{s})-\dot{\varphi}_{s}\right)\right)\,{\rm d}s\right\}$
for all $c\in\mathbb{R}$. We now analyze the term $B_{\cdot 22}$. Under the condition $\|W^{\sigma}\|\leq\epsilon$, we assume without loss of generality that $\mathbb{P}(\|W^{\sigma}\|\leq\epsilon)=\mathbb{P}(\omega_{1})$, meaning that within the probability event $\omega_{1}$, the condition $\|W^{\sigma}\|\leq\epsilon$ holds. Since $\sigma$ satisfies condition $(C2)$, we can conclude that there exists a constant $C$ such that $\|W\|\leq C\epsilon$ holds within the probability event $\omega_{1}$.
We now consider the integrand in $B_{\cdot 22}$, the term $\int_{0}^{s}\left(\sigma(t,y_{t})-\sigma(t,\varphi_{t})\right)\,\mathrm{d}W_{t}$. Based on the Lipschitz continuity of $\sigma$, we have $\sigma(t,y_{t})-\sigma(t,\varphi_{t})\leq L|y_{t}-\varphi_{t}|=L\|W^{\sigma}\|$. Therefore, we can infer that this term is of order $\mathcal{O}(\epsilon^{2})$ within the event $\omega_{1}$. That is, we can write $B_{\cdot 22}$ as
$B_{\cdot 22}=\int_{0}^{1}\left\langle R(s),\,\mathrm{d}W_{s}\right\rangle,$
where $R(s)$ denotes the remainder terms satisfying $\mathbb{E}|R(s)|^{2}=\mathcal{O}(\epsilon^{4})$. Therefore, we will consider $B_{\cdot 22}$ together with $B_{13}$. Last,
$B_{\cdot 22}+B_{13}+B_{23}$ $=\int_{0}^{1}\left\langle{R(s)},\,{\rm d}W_{s}\right\rangle.$
For any $c_{1},c_{2},c_{3}\in\mathbb{R}$, $c=3\max\left\{c_{1},c_{2},c_{3}\right\}$ and $\delta>0$, we have
$\quad\mathbb{E}\left({\rm exp}\left\{c_{1}B_{122}+c_{2}B_{13}+c_{3}B_{23}\right\}\big{|}\|W^{\sigma}\|\leq\epsilon\right)$ (16)
$\leq\int_{0}^{\infty}{e^{x}\mathbb{P}\left(\left|c\int_{0}^{1}\left\langle{R(s)},\,{\rm d}W_{s}\right\rangle\right|>x\big{|}\|W^{\sigma}\|\leq\epsilon\right)}\,{\rm d}x$
$\leq\int_{\delta}^{\infty}{e^{x}\mathbb{P}\left(\left|c\int_{0}^{1}\left\langle{R(s)},\,{\rm d}W_{s}\right\rangle\right|>x\big{|}\|W^{\sigma}\|\leq\epsilon\right)}\,{\rm d}x+e^{\delta}.$
Define the martingale $M(t)=c\int_{0}^{t}\left\langle{R(s)},\,{\rm d}W_{s}\right\rangle$. We have estimate about its quadratic variation
$\langle M(t)\rangle=c^{2}\int_{0}^{t}{\mathbb{E}\left|R(s)\right|^{2}}\,{\rm d}s\leq C\epsilon^{4}$
$\mathbb{P}\left(\left|c\int_{0}^{1}\left\langle{R(s)},\,{\rm d}W_{s}\right\rangle\right|>x,\|W^{\sigma}\|\leq\epsilon\right)\leq{\rm exp}\left\{-\frac{x^{2}}{2C\epsilon^{4}}\right\}.$
Then, by Lemma 2.19, we have
$\mathbb{P}\left(\left|c\int_{0}^{1}\left\langle{R(s)},\,{\rm d}W_{t}\right\rangle\right|>x\big{|}\|W^{\sigma}\|\leq\epsilon\right)\leq{\rm exp}\left\{-\frac{x^{2}}{2C\epsilon^{4}}\right\}{\rm exp}\left\{\frac{C}{\epsilon^{2}}\right\}.$
[…]
$\limsup\limits_{\epsilon\to 0}\mathbb{E}\left({\rm exp}\left\{c_{1}\bar{B}_{22}+c_{2}B_{13}+c_{3}B_{23}\right\}\big{|}\|W^{\sigma}\|\leq\epsilon\right)=1$ (17)
for all $c_{1},c_{2},c_{3}\in\mathbb{R}$, as $\epsilon\to 0$ and $\delta\to 0$.
For the third term $B_{3}$, let $v_{t}:=J\nabla H(y_{t})-\dot{\varphi}_{t}$, then
$B_{3}$ $=\frac{1}{2}\int_{0}^{1}\left(\left|\sigma^{-1}(s,\varphi_{s})v_{s}\right|^{2}-\left|\sigma^{-1}(s,y_{s})v_{s}\right|^{2}\right)\,{\rm d}s$
$=\frac{1}{2}\int_{0}^{1}\left\langle\left(\left({\sigma^{-1}\left(s,y_{s}\right)}\right)^{2}-\left({\sigma^{-1}\left(s,\varphi_{s}\right)}\right)^{2}\right)v_{s},v_{s}\right\rangle\,{\rm d}s$
$\leq\frac{1}{2}\int_{0}^{1}{\left(\left\|\left({\sigma^{-1}\left(s,y_{s}\right)}\right)^{2}-\left({\sigma^{-1}\left(s,\varphi_{s}\right)}\right)^{2}\right\|_{2}\right)\left|v_{s}\right|^{2}}\,{\rm d}s$
$\leq\frac{1}{2}\int_{0}^{1}{\left\|{\sigma^{-1}\left(s,y_{s}\right)}-{\sigma^{-1}\left(s,\varphi_{s}\right)}\right\|_{2}\left\|{\sigma^{-1}\left(s,y_{s}\right)}+{\sigma^{-1}\left(s,\varphi_{s}\right)}\right\|_{2}\left|v_{s}\right|^{2}}\,{\rm d}s$
$\leq\frac{1}{2}\int_{0}^{1}\left\|{\sigma^{-1}\left(s,y_{s}\right)}\right\|_{2}\left\|{\sigma\left(s,\varphi_{s}\right)-\sigma\left(s,y_{s}\right)}\right\|_{2}\left\|{\sigma^{-1}\left(s,\varphi_{s}\right)}\right\|_{2}$
$\qquad\times\left\|{\sigma^{-1}\left(s,y_{s}\right)}+{\sigma^{-1}\left(s,\varphi_{s}\right)}\right\|_{2}\left|v_{s}\right|^{2}\,{\rm d}s.$
Since both $\sigma^{-1}$ and $v$ are bounded, and $\sigma$ is Lipschitz continuous, while $\varphi_{t}-X_{0}\in\mathbb{H}^{1}$, we conclude that there exists a constant $C>0$ such that
$B_{3}\leq C\int_{0}^{1}\left\|\sigma(s,\varphi_{s})-\sigma(s,y_{s})\right\|_{2}\,{\rm d}s\leq CL\left\|W^{\sigma}\right\|.$
The above inequality implies that
$\limsup\limits_{\epsilon\to 0}\mathbb{E}\left({\rm exp}\left\{cB_{3}\right\}\big{|}\|W^{\sigma}\|\leq\epsilon\right)=1$ (18)
@@
For the fourth term $B_{4}$, let $v_{t}:=J\nabla H(y_{t})-\dot{\varphi}_{t}$, $u:=J\nabla H(\varphi_{t})-\dot{\varphi}_{t}$, then
$B_{4}$ $=\frac{1}{2}\int_{0}^{1}{\left|\sigma^{-1}(s,\varphi_{s})u\right|^{2}-\left|\sigma^{-1}(s,\varphi_{s})v\right|^{2}}\,{\rm d}s$
$=\frac{1}{2}\int_{0}^{1}{\left\langle\sigma^{-1}(s,\varphi_{s})\left(u-v\right),\sigma^{-1}(s,\varphi_{s})\left(u+v\right)\right\rangle}\,{\rm d}s$
$=\frac{1}{2}\int_{0}^{1}{\left\langle\sigma^{-1}(s,\varphi_{s})^{2}\left(u-v\right),\left(u+v\right)\right\rangle}\,{\rm d}s$
$\leq\frac{1}{2}\int_{0}^{1}{\left\|\sigma^{-1}(s,\varphi_{s})^{2}\right\|_{2}\left\|\left(u-v\right)\right\|\left\|\left(u+v\right)\right\|}\,{\rm d}s.$
Since both $\sigma^{-1}$, $v$, and $u$ are bounded, and $J\nabla H(y_{t})$ is Lipschitz continuous, and noting that $\varphi_{t}-x_{0}\in\mathbb{H}^{1}$, we can conclude that there exists a constant $C>0$ such that
$B_{4}\leq C\left|u-v\right|\leq C\int_{0}^{1}\left|J\nabla H(\varphi_{s})-J\nabla H(y_{s})\right|\,{\rm d}s\leq CL\left\|W^{\sigma}\right\|.$
So we have
$\limsup\limits_{\epsilon\to 0}\mathbb{E}\left({\rm exp}\left\{cB_{4}\right\}\big{|}\|W^{\sigma}\|\leq\epsilon\right)=1$ (19)
@@
$\lim\limits_{\epsilon\to 0}\frac{\mathbb{P}\left(\|X_{t}-\varphi_{t}\|\leq\epsilon\right)}{\mathbb{P}\left(\|W\|\leq\epsilon\right)}$ $={\rm exp}\left\{-\frac{1}{2}\int_{0}^{1}{\left|\sigma^{-1}(t,\varphi_{s})\left(J\nabla H(\varphi_{s})-\dot{\varphi_{s}}\right)\right|^{2}}\,{\rm d}s\right$.
@@
$\qquad\quad\left.+\frac{1}{2}\int_{0}^{1}\mathrm{Tr}\left(\left(D_{y}\sigma(s,\varphi_{s})\right)\sigma^{-1}(s,\varphi_{s})\left(J\nabla H(\varphi_{s})-\dot{\varphi}_{s}\right)\right)\,{\rm d}s\right\},$
where the term $D_{y}\sigma(t,\varphi_{t})$ refers to the third-order $n\times n\times n$ derivative tensor of the matrix-valued function $\sigma(t,y_{t})$ with respect to $y_{t}\in\mathbb{R}^{n}$, evaluated at $\varphi_{t}$. Consequently,
$OM(\varphi,\dot{\varphi})=\int_{0}^{1}{\left|\sigma^{-1}(t,\varphi_{s})\left(J\nabla H(\varphi_{s})-\dot{\varphi_{s}}\right)\right|^{2}}\,{\rm d}s-\int_{0}^{1}\mathrm{Tr}\left(\left(D_{y}\sigma(s,\varphi_{s})\right)\sigma^{-1}(s,\varphi_{s})\left(J\nabla H(\varphi_{s})-\dot{\varphi}_{s}\right)\right)\,{\rm d}s.$
4 Most Probable Path
In this section, building upon Theorem 3.1, we further introduce the method of Euler–Lagrange equations to study the most probable paths in the solution trajectories of stochastic differential equations. It is well known that, under small noise limits or small-ball probability constraints, the maximum likelihood path of a stochastic process tends to minimize the so-called Onsager-Machlup functional, given by
$OM(\varphi,\dot{\varphi})=\int_{0}^{1}{\left|\sigma^{-1}(t,\varphi_{s})\left(J\nabla H(\varphi_{s})-\dot{\varphi_{s}}\right)\right|^{2}}\,{\rm d}s-\int_{0}^{1}\mathrm{Tr}\left(\left(D_{y}\sigma(s,\varphi_{s})\right)\sigma^{-1}(s,\varphi_{s})\left(J\nabla H(\varphi_{s})-\dot{\varphi}_{s}\right)\right)\,{\rm d}s.$
The integrand
$\left|\sigma^{-1}(t,\varphi_{s})\left(J\nabla H(\varphi_{s})-\dot{\varphi_{s}}\right)\right|^{2}-\mathrm{Tr}\left(\left(D_{y}\sigma(s,\varphi_{s})\right)\sigma^{-1}(s,\varphi_{s})\left(J\nabla H(\varphi_{s})-\dot{\varphi}_{s}\right)\right)$
can be viewed as the Lagrangian function associated with this variational problem. According to the classical calculus of variations, the optimal path $\varphi$ should satisfy the corresponding Euler–Lagrange equation
$\frac{\partial L}{\partial\varphi}-\frac{\mathrm{d}}{\mathrm{d}t}\left(\frac{\partial L}{\partial\dot{\varphi}}\right)=0,$
together with the initial condition $\varphi_{0}=X_{0}$ and an appropriate terminal condition.
Due to the presence of nonlinear matrix-valued functions, trace operators, and nontrivial drift structures, the above equation is typically challenging to solve explicitly. However, it plays a critical role in both theoretical analysis and numerical investigations. To address this challenge, we adopt the Ansatz method as a more tractable and analytically convenient approach. Accordingly, we present the following theorem on the solution of the maximum probability path.
Theorem 4.1.